OpenStudy (anonymous):

Determine which of the following expressions represent the exact differential and find the corresponding function satisfying the condition f(2,-1) = 2 - 5ln2 if such function exists. a) (sin^2(xy/2) + e^x/y) dx - (cos(2x/y) - e^x/y) dy; b) (1/y^2-5/x) dx + 2/y(5-(x/y^2) dy

OpenStudy (amistre64):

an exact diff. is one where you \(\int M(x,y) dx\) and add on a function in y, then \(\frac{d}{dx}\) down into N(x,y) i think

OpenStudy (amistre64):

or vice versa

OpenStudy (amistre64):

\[(sin^2(\frac{xy}{2}) + e^{\frac{x}{y}}) dx - (cos(\frac{2x}{y}) - e^{\frac{x}{y}}) dy\] \[\int sin^2(\frac{xy}{2}) + e^{\frac{x}{y}} dx\] \[\int sin^2(\frac{xy}{2})dx + ye^{\frac{x}{y}} \] and what is the reduction formula for sins?

OpenStudy (amistre64):

sin^2(x) = 1/2 -cos(2x)/2 i think

OpenStudy (amistre64):

\[\int (\frac{1}{2}-\frac{1}{2}cos(2\frac{xy}{2}))dx + ye^{\frac{x}{y}}\] \[\frac{1}{2}x-\frac{1}{2}y\ sin(xy) + ye^{\frac{x}{y}}\]

OpenStudy (amistre64):

err... 1/2y on the sin

OpenStudy (amistre64):

\[\frac{d}{dy}\ \frac{1}{2}x-\frac{1}{2y}\ sin(xy) + ye^{\frac{x}{y}}+g(y)\] \[\frac{d}{dy}(-\frac{1}{2y}\ sin(xy) + ye^{\frac{x}{y}})+g'(y)\]

OpenStudy (amistre64):

\[\frac{d}{dy}(-\frac{1}{2y}\ sin(xy))=-\frac{1}{2y}'\ sin(xy)+\frac{-1}{2y}\ sin(xy)' \] \[=>\ \frac{1}{2y^2}\ sin(xy)-\frac{1}{2y}\ x\ cos(xy) \] \[\frac{d}{dy}ye^{\frac{x}{y}}=y'e^{\frac{x}{y}}+ye'^{\frac{x}{y}}\] \[=>\ e^{\frac{x}{y}}+y \frac{x}{y}'e^{\frac{x}{y}}\] \[=>\ e^{\frac{x}{y}}+y \frac{-x}{y^2}e^{\frac{x}{y}}\] \[=>\ e^{\frac{x}{y}}-\frac{x}{y}e^{\frac{x}{y}}\] \[\frac{1}{2y^2}\ sin(xy)-\frac{1}{2y}\ x\ cos(xy) + e^{\frac{x}{y}}-\frac{x}{y}e^{\frac{x}{y}}+g'(y)\] maybe?

OpenStudy (amistre64):

and equate that to the dy side

OpenStudy (amistre64):

\[\small\frac{1}{2y^2}\ sin(xy)-\frac{1}{2y}\ x\ cos(xy) + e^{\frac{x}{y}}-\frac{x}{y}e^{\frac{x}{y}}+g'(y)=cos(\frac{2x}{y}) - e^{\frac{x}{y}}\] what a mess ....

OpenStudy (amistre64):

\[\small\frac{sin(xy)-xy\ cos(xy)-2xy\ e^{\frac{x}{y}}+2y^2 e^{\frac{x}{y}}}{2y^2}+g'(y)=\frac{2y^2cos(\frac{2x}{y}) - 2y^2e^{\frac{x}{y}}}{2y^2}\] \[\small{sin(xy)-xy\ cos(xy)-2xy\ e^{\frac{x}{y}}+2y^2 e^{\frac{x}{y}}}+g'(y)={2y^2cos(\frac{2x}{y}) - 2y^2e^{\frac{x}{y}}}\] \[\small g'(y)={-sin(xy)+xycos(xy)+2xye^{\frac{x}{y}}-2y^2 e^{\frac{x}{y}}}+{2y^2cos(\frac{2x}{y})-2y^2e^{\frac{x}{y}}}\] then we either simplify and up that mess or we up it and let the undergrads sort thru it

OpenStudy (anonymous):

Thanks for your help would you know what I need to do to complete the question correctly? This is for an assignment for university? I am stuck.

OpenStudy (amistre64):

i would have to rework it to make sure i didnt mess any of it up

OpenStudy (anonymous):

ok only if you have time. would be appreciated final maths course and i want to pass it.

OpenStudy (amistre64):

so far its correct down to the "maybe" part :) \[\frac{-2 y e^{\frac{x}{y}} (x-y)+sin(x y)-xy\ cos(x y)}{2 y^2}+g'(y)\] the missing part; the g'(y) is what should make this equal to what is given already in the "dy" part of the problem.

OpenStudy (amistre64):

\["that"+g'(y)=cos(2x/y) - e^{x/y}\] \[g'(y)=cos(\frac{2x}{y}) - e^{x/y}-"that"\] \[\small g'(y)=cos(\frac{2x}{y}) - e^{x/y}-\left(\frac{-2 y e^{\frac{x}{y}} (x-y)+sin(x y)-xy\ cos(x y)}{2 y^2}\right)\] \[ g'(y)=cos(\frac{2x}{y}) - e^{x/y}-\frac{2 y e^{\frac{x}{y}} (x-y)-sin(x y)+xy\ cos(x y)}{2 y^2}\]

OpenStudy (amistre64):

i spose it would be easier to split it up into parts and work its integration

OpenStudy (amistre64):

unless there is an easier way that i cant recall, this thing turns out ugly

OpenStudy (amistre64):

we could try to up the dy and bring it down to the dx and see if that is a nicer fit

OpenStudy (jamesj):

No, it's easier than all this. A differential P dx + Q dy is exact if thereis a function F such that dF = P dx + Q dy I.e., \[P = \partial F / \partial x \ \ \text{ and } \ \ Q = \partial F/\partial y\] So for the first example, (sin^2(xy/2) + e^x/y) dx - (cos(2x/y) - e^x/y) dy it is clear that there is no such F because just looking at hte e^x/y terms, there is no function G such that that partial deriv wrt x gives e^x/y and the partial deriv wrt y gives -e^x/y Formalize that by using the two pdes above if you like. Now for the second example, isn't written correctly. Can you give it to us again?

OpenStudy (amistre64):

yeah, those y denominators in the dy part was causing me concern .... when i tried to correlate it with the wolf, it spit back it "ugly" answers ...

OpenStudy (anonymous):

(1/y^2-5/x) dx + 2/y(5-x/y^2)dy